# On the Training Convergence of Transformers for In-Context Classification

### v1

@@ Abstract
While transformers have demonstrated impressive capacities for in-context learning (ICL) in practice, theoretical understanding of the underlying mechanism enabling transformers to perform ICL is still in its infant stage. This work aims to theoretically study the training dynamics of transformers for in-context classification tasks. We demonstrate that, for in-context classification of Gaussian mixtures under certain assumptions, a single-layer transformer trained via gradient descent converges to a globally optimal model at a linear rate. We further quantify the impact of the training and testing prompt lengths on the ICL inference error of the trained transformer. We show that when the lengths of training and testing prompts are sufficiently large, the prediction of the trained transformer approaches the Bayes-optimal classifier. Experimental results corroborate the theoretical findings.
@@ 1 Introduction
This impressive capacity of transformer-based models has inspired many recent works aiming to understand the ICL abilities of transformers. A more comprehensive literature review can be found in Appendix A. Garg et al. (2022) was the first to study the ICL abilities of transformers for various function classes. They empirically showed that transformers can learn linear regression models in context. Later on, a line of research was developed to theoretically explain how transformers perform in-context linear regression. For example, Akyürek et al. (2022); Von Oswald et al. (2023); Bai et al. (2024); Fu et al. (2023); Giannou et al. (2024) showed by construction that, some specially-designed transformers can perform linear regression in context. Moreover, some recent works like Zhang et al. (2023a); Huang et al. (2023); Chen et al. (2024) studied the training dynamics of a single-layer transformer for in-context linear regression. They proved the convergence of their single-layer transformers and showed their trained transformer are able to perform linear regression in context.
@@
Building on the earlier works that largely focus on linear regression problems, several recent papers have started to investigate the ICL capabilities of transformers for non-linear problems such as classification. For instance, Bai et al. (2024) showed that, by construction, multi-layer transformers can be approximately viewed as multiple steps of gradient descents for logistic regression. Giannou et al. (2024) further showcased that the constructed transformers can approximately perform Newton’s method for logistic regression. Recently, Li et al. (2024) studied the training dynamics of transformers for in-context binary classification. However, their analysis requires the data to be pairwise orthogonal and the possible distribution of their data is highly limited. The learning dynamics of transformers for more general in-context classification problems is not well understood. Moreover, to the best of our knowledge, existing literature (Bai et al., 2024; Giannou et al., 2024; Li et al., 2024) studying the in-context classification of transformers focus only on binary classification. How transformers perform in-context multi-class classification remains unexplored.
@@
In this work, we study the learning dynamics of a singly-layer transformer for both in-context binary and multi-class classification of Gaussian mixtures, a fundamental problem in machine learning. Our main contributions can be summarized as follows:
@@
  • To the best of our knowledge, we are the first to study the learning dynamics of transformers for in-context classification of Gaussian mixtures, and we are the first to prove the training convergence of transformers for in-context multi-class classification. We prove that with appropriately distributed training data (Assumptions 3.1, 4.1), a single-layer transformer trained via gradient descent will converge to its global minimizer at a linear rate (Theorems 3.1, 4.1) for both in-context binary or multi-class classification problems.
@@
  • Due to the high non-linearity of our loss function, we cannot directly find the closed-form expression of the global minimizer. Instead, we prove an important property that the global minimizer consists of a constant plus an error term that is induced by the finite training prompt length ($N$). We further show that the max norm of this error term is bounded, and converges to zero at a rate of $O(1/N)$.
@@
  • With properly distributed test prompts (Assumptions 3.2, 4.2), we establish an upper bound of the inference error (defined in Equation (3)) of the trained transformer and quantify the impact of the training and testing prompt lengths on this error. We further prove that when the lengths of training prompts ($N$) and testing prompts ($M$) approach infinity, this error converges to zero at a rate of $O(1/N+1/\sqrt{M})$ (Theorems 3.2, 4.2), and the prediction of the trained transformer is Bayes-optimal, i.e., the optimal classifier given the data distribution.
@@ 2.1 Single-layer transformer
Given an input embedding matrix $E\in\mathbb{R}^{d_{e}\times d_{n}}$, a single head self-attention module $F_{SA}$ (Vaswani et al., 2017) with width $d_{e}$ will output
$F_{SA}(E;W^{V},W^{K},W^{Q})=E+W^{V}E\cdot f_{\text{attn}}\left(\frac{(W^{K}E)^{\top}W^{Q}E}{\rho}\right),$ (1)
@@
In the following theoretical study and the subsequent experiments (Section 5.2), we show that this simplified transformer model has sufficient capability to approach the Bayes-optimal classifier for in-context classification of Gaussian mixtures.
@@ 2.2 In-context learning framework
$\Delta(y_{\mathsf{query}},\widehat{y}_{\mathsf{query}})=\sup_{z\in R(y_{\mathsf{query}})}|\mathbb{P}\left(y_{\mathsf{query}}=z\right)-\mathbb{P}\left(\widehat{y}_{\mathsf{query}}=z\right)|,$ (3)
@@
Unlike standard supervised learning, in ICL, each prompt $P_{\tau}$ can be sampled from a different distribution $\mathcal{P}_{\tau}$. We say that a model has the ICL capability if it can approximate $y_{\tau,\mathsf{query}}$ for a broad range of $\mathcal{P}_{\tau}$’s with fixed parameters.
@@ 3 In-context binary classification
$E=E(P)=\begin{pmatrix}x_{1}&x_{2}&\cdots&x_{N}&x_{\mathsf{query}}\\ y_{1}&y_{2}&\cdots&y_{N}&0\end{pmatrix}\in\mathbb{R}^{(d+1)\times(N+1)}.$ (4)
@@
$=\sigma\Bigg{(}\begin{pmatrix}0_{d}^{\top}&1\end{pmatrix}\begin{pmatrix}\frac{1}{N}\sum_{i=1}^{N}x_{i}x_{i}^{\top}+\frac{1}{N}x_{\mathsf{query}}x_{\mathsf{query}}^{\top}&\frac{1}{N}\sum_{i=1}^{N}x_{i}y_{i}\\ \frac{1}{N}\sum_{i=1}^{N}x_{i}^{\top}y_{i}&\frac{1}{N}\sum_{i=1}^{N}y_{i}^{2}\end{pmatrix}\begin{pmatrix}W&0_{d}\\ 0_{d}^{\top}&0\end{pmatrix}\begin{pmatrix}x_{\mathsf{query}}\\ 0\end{pmatrix}\Bigg{)}$
@@ 3.1 Training procedure
    $\{x_{\tau,i},y_{\tau,i}\}_{i=1}^{N},\{x_{\tau,\mathsf{query}},y_{\tau,\mathsf{query}}\}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{P}^{b}(\mu_{\tau,0},\mu_{\tau,1},\Lambda).$
@@
    $\mu_{\tau,0}$ is randomly sampled from $\mathsf{N}(0,I_{d})$, and $\mu_{\tau,1}=U_{\tau,\Lambda}\mu_{\tau,0}$ where $U_{\tau,\Lambda}=\Lambda^{1/2}U_{\tau}\Lambda^{-1/2}$, and $U_{\tau}$ is uniformly distributed over the closed set of real unitary matrices such that $U_{\tau}U_{\tau}^{\top}=I_{d}$.
@@
$\widehat{L}(W)=\frac{1}{2B}\sum_{\tau=1}^{B}-(1+y_{\tau,\mathsf{query}})\log(\widehat{y}_{\tau,\mathsf{out}})-(1-y_{\tau,\mathsf{query}})\log(1-\widehat{y}_{\tau,\mathsf{out}}).$ (6)
@@
$L(W)=\lim_{B\to\infty}\widehat{L}(W)=-\frac{1}{2}\mathbb{E}\left[(1+y_{\tau,\mathsf{query}})\log(\widehat{y}_{\tau,\mathsf{out}})+(1-y_{\tau,\mathsf{query}})\log(1-\widehat{y}_{\tau,\mathsf{out}})\right],$ (7)
@@
    Optimizing the training loss $L(W)$ in (7) with training prompt length $N$ via gradient descent $W^{t+1}=W^{t}-\eta\nabla L(W^{t})$, we have that for any $t\geq 1$
@@
    Denote $W^{*}=2(\Lambda^{-1}+G)$, $q=x_{\tau,\mathsf{query}}$, $\mu=\mu_{\tau,1}-\mu_{\tau,0}$, $u=2(\mu_{\tau,1}+\mu_{\tau,0})$, $a=\mu^{\top}\Lambda^{-1}q$ for simplicity. Then we have
    $\|G\|_{\max}\leq$ $\frac{1}{N}\Big{\|}S^{-1}\mathbb{E}\big{[}\sigma^{\prime}(a)\left(4qq^{\top}+\frac{1}{4}uu^{\top}\Lambda^{-1}qq^{\top}\right)$
    $+\sigma^{\prime\prime}(a)\left(\frac{1}{8}(u^{\top}\Lambda^{-1}q)^{2}\mu q^{\top}+2q^{\top}\Lambda^{-1}q\mu q^{\top})\big{]}\right)\Big{\|}_{\max}+o\left(\frac{1}{N}\right),$ (10)
@@
Proof sketch for Theorem 3.1. As a first step, we prove in Lemma D.2 that the expected loss function $L(W)$ in (7) is strictly convex with respect to (w.r.t.) $W$ and is strongly convex in any compact set of $\mathbb{R}^{d\times d}$. Moreover, we prove $L(W)$ has one unique global minimizer $W^{*}$. Since the loss function $L(W)$ we consider is highly non-linear, we cannot directly find the closed-form expression of $W^{*}$, as is often done in the prior literature. We address this technical challenge via the following method. First, in Lemma D.3, by analyzing the Taylor expansion of $L(W)$, we prove that as $N\to\infty$, our loss function $L(W)$ converges to $\widetilde{L}(W)$ pointwisely (defined in (25)), and the global minimizer $W^{*}$ converges to $2\Lambda^{-1}$. Thus, we denote $W^{*}=2(\Lambda^{-1}+G)$, and prove $\|G\|_{\max}$ is bounded and scales as $\|G\|_{\max}=O(N^{-1/2})$. Next, in Lemma D.4, by further analyzing the Taylor expansion of the equation $\nabla L(W^{*})=0$ at the point $2\Lambda^{-1}$, we establish a tighter bound $\|G\|_{\max}=O(N^{-1})$. In Lemma D.5, we prove that our loss function is $l$-smooth and provide an upper bound for $l$. Thus, in a compact set $R_{W}$, our loss function is $\alpha$-strongly convex and $l$-smooth. Finally, leveraging the standard results from the convex optimization, we prove Theorem 3.1.
@@
Corollary 3.1
If we optimize the expected loss $L(W)$ in (7) via gradient descent with training prompt length $N$, initial parameters $W^{0}$, and learning rate $\eta=1/l$, then, under Assumption 3.1, after $T\geq 2\kappa\log(N\|W^{0}-W^{*}\|_{F})$ steps, the updated model $\widehat{W}$ satisfies
@@ 3.2 In-context inference
Next, we analyze the performance of the trained transformer (11) for in-context binary classification tasks. We make the following assumption.
Assumption 3.2
@@
$\mathbb{P}\left(y_{\mathsf{query}}=1|x_{\mathsf{query}}\right)=$ $\frac{f(x_{\mathsf{query}}|y_{\mathsf{query}}=1)\mathbb{P}\left(y_{\mathsf{query}}=1\right)}{\sum_{z\in\{\pm 1\}}f(x_{\mathsf{query}}|y_{\mathsf{query}}=z)\mathbb{P}\left(y_{\mathsf{query}}=z\right)}=\sigma((\mu_{1}-\mu_{0})^{\top}\Lambda^{-1}x_{\mathsf{query}}).$
If we test the trained transformer with parameters $\widehat{W}$ in (11) and $P_{\mathsf{test}}$, by a simple calculation, we have
@@
Theorem 3.2
Consider a test prompt $P_{\mathsf{test}}$ satisfying Assumption 3.2, and let $y_{\mathsf{query}}\sim\mathcal{P}^{b}_{y|x_{\mathsf{query}}}(\mu_{0},\mu_{1},\Lambda)$. Let $\widehat{y}_{\mathsf{query}}$ be the prediction of the trained transformer with parameters $\widehat{W}$ in (11). Then, for the inference error defined in (3), we have
@@
$\leq\sigma^{\prime}(\mu^{\top}\Lambda^{-1}q)\left[\|\widehat{G}\|_{\max}\sum_{i,j\in[d]}|\mu_{i}q_{j}|+\frac{1}{\sqrt{M}}\left(\frac{1}{2}|u^{\top}\Lambda^{-1}q|+\frac{2\sqrt{2}}{\sqrt{\pi}}\sum_{i,j\in[d]}|\Lambda^{-1/2}_{ij}q_{j}|\right)\right]+o\left(\frac{1}{N}+\frac{1}{\sqrt{M}}\right),$
@@
The proof of Theorem 3.2 can be found in Appendix E. Since $\|\widehat{G}\|_{\max}=O(1/N)$, Theorem 3.2 suggests that if we ignore the constants regarding $\mu_{0},\mu_{1},\Lambda,x_{\mathsf{query}}$, the expected total variation distance between $y_{\mathsf{query}}$ and $\widehat{y}_{\mathsf{query}}$ is at most $O(1/N+1/\sqrt{M})$. On the other hand, for data pair $(x,y)\sim\mathcal{P}^{b}(\mu_{0},\mu_{1},\Lambda)$, the Bayes-optimal classifier is $\mathbb{P}(y=1|x)=f(x|y)\mathbb{P}(y=1)/f(x)=\sigma((\mu_{1}-\mu_{0})^{\top}\Lambda^{-1}x)$, which corresponds to the logistic regression model $\sigma(w^{\top}x+b)$ with parameters $w=\Lambda^{-1}(\mu_{1}-\mu_{0})$ and $b=0$. Therefore, when $N,M\to\infty$, the prediction of the trained transformer is Bayes-optimal, and is equivalent to the optimal logistic regressor for binary classification problems with distribution $\mathcal{P}^{b}(\mu_{0},\mu_{1},\Lambda)$. Note that different from Assumption 3.1 which states that $\mu_{\tau,0},\mu_{\tau,1},x_{\tau,\mathsf{query}}$ are sampled according to some specific distributions during training, Assumption 3.2 does not impose strong distributional constraints on $\mu_{0},\mu_{1}$ and $x_{\mathsf{query}}$, which shows the strong generalization ability of the trained transformer. We also discuss the consequences when Assumption 3.2 does not hold in Remark E.1, which highlights the necessity of Assumption 3.2. Moreover, even if $M\to\infty$, the distribution variation between $y_{\mathsf{query}}$ and $\widehat{y}_{\mathsf{query}}$ does not disappear unless $N\to\infty$. Thus, the ICL ability of trained transformers for binary classification is limited by the finite length of training prompts. Similar behaviors have also been observed in Zhang et al. (2023a) for in-context linear regression.

### v3

@@ Abstract
Although transformers have demonstrated impressive capabilities for in-context learning (ICL) in practice, theoretical understanding of the underlying mechanism that allows transformers to perform ICL is still in its infancy. This work aims to theoretically study the training dynamics of transformers for in-context classification tasks. We demonstrate that, for in-context classification of Gaussian mixtures under certain assumptions, a single-layer transformer trained via gradient descent converges to a globally optimal model at a linear rate. We further quantify the impact of the training and testing prompt lengths on the ICL inference error of the trained transformer. We show that when the lengths of training and testing prompts are sufficiently large, the prediction of the trained transformer approaches the ground truth distribution of the labels. Experimental results corroborate the theoretical findings.
Machine Learning, ICML
@@ 1 Introduction
This impressive capacity of transformer-based models has inspired many recent works aiming to understand the ICL abilities of transformers. A more comprehensive literature review can be found in Appendix A. Garg et al. (2022) was the first to study the ICL abilities of transformers for various function classes. They empirically showed that transformers can learn linear regression models in context. Later on, a line of research was developed to theoretically explain how transformers perform in-context linear regression. For example, Akyürek et al. (2022); Von Oswald et al. (2023); Bai et al. (2024); Fu et al. (2023); Giannou et al. (2024) showed by construction that, some specially-designed transformers can perform linear regression in context. Moreover, some recent works such as Zhang et al. (2023a); Huang et al. (2023); Chen et al. (2024) studied the training dynamics of a single-layer transformer for in-context linear regression. They proved the convergence of certain single-layer transformers during training and showed that the trained transformers are able to perform linear regression in context.
@@
Building on the earlier works that largely focus on linear regression problems, several recent papers have started to investigate the ICL capabilities of transformers for classification problems. For instance, Bai et al. (2024) showed that, by construction, multi-layer transformers can be approximately viewed as multiple steps of gradient descents for logistic regression. Giannou et al. (2024) further showcased that the constructed transformers can approximately perform Newton’s method for logistic regression. Lin & Lee (2024) studied the dual operating modes for in-context classification of Gaussian mixtures. However, their analyses were based on the idealized Bayes-optimal next-token predictor for ICL tasks and did not consider the training dynamics of transformer models. Some recent works have started to study the training dynamics of transformers for certain classification problems. For example, Li et al. (2024b) proved that a single-layer transformer can be trained to learn one-nearest neighbor in context. Li et al. (2024a) studied the training dynamics of a single-layer transformer for some binary classification tasks with finite, pairwise orthogonal patterns. Frei & Vardi (2024) analyzed a linear transformer model for in-context classification of Gaussian mixtures, assuming additional conditions such as a sufficiently large signal-to-noise ratio. However, all these works (Li et al., 2024a, b; Frei & Vardi, 2024) only considered the binary classification problems. The training dynamics of transformers for more general in-context classification problems beyond the specific settings and assumptions in Li et al. (2024a, b); Frei & Vardi (2024) remain largely under-explored.
@@
In this work, we study the training dynamics of a single-layer transformer for both binary and multi-class classification of Gaussian mixtures during ICL, a fundamental problem in machine learning. Our main contributions can be summarized as follows:
@@
  • We prove that with appropriately distributed training data (Assumptions 3.2, 4.2), a single-layer transformer trained via gradient descent will converge to its global minimizer at a linear rate (Theorems 3.3, 4.3) for both in-context binary or multi-class classification problems. To the best of our knowledge, we are the first to prove the training convergence of transformers for in-context multi-class classification. Moreover, our analysis reveals that the trained single-layer transformer can be viewed as approximately implementing linear discriminant analysis (LDA).
@@
  • Due to the non-linearity of our loss function, we cannot directly find the closed-form expression of the global minimizer. Instead, we prove an important property that the global minimizer consists of a constant plus an error term that is induced by the finite training prompt length ($N$). We further show that the max norm of this error term is bounded, and converges to zero at a rate of $O(1/N)$.
@@
  • With properly distributed testing prompts (Assumptions 3.5, 4.4), we establish an upper bound of the inference error (defined in Equation (3)) of the trained transformer and quantify the impact of the training and testing prompt lengths on this error. We further prove that when the lengths of training prompts ($N$) and testing prompts ($M$) approach infinity, this error converges to zero at a rate of $O(1/N+1/\sqrt{M})$ , and the prediction of the trained transformer has an identical distribution to that of the ground-truth label (Theorems 3.6, 4.5).
@@ 2.1 Single-layer transformer
Given an input embedding matrix $E\in\mathbb{R}^{d_{e}\times d_{n}}$, a single head self-attention module $F_{SA}$ with width $d_{e}$ will output
$F_{SA}(E;W^{V},W^{K},W^{Q})$
$=$ $E+W^{V}E\cdot f_{\text{attn}}\left(\frac{(W^{K}E)^{\top}W^{Q}E}{\rho}\right),$ (1)
@@
In the following theoretical study and the subsequent experiments (Section 5.2), we show that this simplified transformer model has sufficient capability to approach the optimal classifier for the in-context classification of Gaussian mixtures.
@@ 2.2 In-context learning framework
$\Delta(y_{\mathsf{query}},\widehat{y}_{\mathsf{query}})$
$=$ $\sup_{z\in R(y_{\mathsf{query}})}|\mathbb{P}\left(y_{\mathsf{query}}=z\right)-\mathbb{P}\left(\widehat{y}_{\mathsf{query}}=z\right)|,$ (3)
@@
Unlike standard supervised learning, each prompt $P_{\tau}$ can be sampled from a different distribution $\mathcal{P}_{\tau}$ in ICL. We say that a model has the ICL capability if it can approximate $y_{\tau,\mathsf{query}}$ for a broad range of $\mathcal{P}_{\tau}$’s with fixed parameters.
@@ 3 In-context binary classification
$E=E(P)=\begin{pmatrix}x_{1}&x_{2}&\cdots&x_{N}&x_{\mathsf{query}}\\ y_{1}&y_{2}&\cdots&y_{N}&0\end{pmatrix}.$ (4)
@@ 3.1 Training procedure
    $\{x_{\tau,i},y_{\tau,i}\}_{i=1}^{N}$ and $\{x_{\tau,\mathsf{query}},y_{\tau,\mathsf{query}}\}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{P}^{b}(\mu_{\tau,0},\mu_{\tau,1},\Lambda)$.
@@
    $\mu_{\tau,0}$ is randomly sampled from $\mathsf{N}(0,I_{d})$, $\mu_{\tau,1}=U_{\tau,\Lambda}\mu_{\tau,0}$ where $U_{\tau,\Lambda}=\Lambda^{1/2}U_{\tau}\Lambda^{-1/2}$, and $U_{\tau}$ is uniformly distributed over the closed set of real unitary matrices such that $U_{\tau}U_{\tau}^{\top}=I_{d}$.
@@
$\widehat{L}(W)=$ $\frac{1}{2B}\sum_{\tau=1}^{B}-(1+y_{\tau,\mathsf{query}})\log(\widehat{y}_{\tau,\mathsf{out}})$
$-(1-y_{\tau,\mathsf{query}})\log(1-\widehat{y}_{\tau,\mathsf{out}}).$ (6)
@@
$L(W)=$ $\lim_{B\to\infty}\widehat{L}(W)=-\frac{1}{2}\mathbb{E}[(1+y_{\tau,\mathsf{query}})\log(\widehat{y}_{\tau,\mathsf{out}})$
$+(1-y_{\tau,\mathsf{query}})\log(1-\widehat{y}_{\tau,\mathsf{out}})],$ (7)
@@
    Optimizing the training loss $L(W)$ in Equation (7) with training prompt length $N$ via gradient descent $W^{t+1}=W^{t}-\eta\nabla L(W^{t})$, we have that for any $t\geq 1$,
@@
    $\alpha\leq\lambda_{\min}(\nabla^{2}L(W))\leq\lambda_{\max}(\nabla^{2}L(W))\leq l,$ (9)
    where $\alpha>0$, $l<\infty$, $W\in R_{W},R_{W}=\{W\in\mathbb{R}^{d\times d}\,|\,\|W-W^{*}\|_{F}\leq\|W^{0}-W^{*}\|_{F}\}$.
@@
    Define $G=\frac{1}{2}W^{*}-\Lambda^{-1}$, $q=x_{\tau,\mathsf{query}}$, $\mu=\mu_{\tau,1}-\mu_{\tau,0}$, $u=2(\mu_{\tau,1}+\mu_{\tau,0})$, and $a=\mu^{\top}\Lambda^{-1}q$ for simplicity. Then we have
    $\|G\|_{\max}$
    $\leq$ $\frac{1}{N}\|S^{-1}(\mathbb{E}[\sigma^{\prime}(a)(4qq^{\top}+\frac{1}{4}uu^{\top}\Lambda^{-1}qq^{\top})$
    $+\sigma^{\prime\prime}(a)(\frac{1}{8}(u^{\top}\Lambda^{-1}q)^{2}\mu q^{\top}+2q^{\top}\Lambda^{-1}q\mu q^{\top})])\|_{\max}$
    $+o(1/N),$ (10)
@@
Proof sketch for Theorem 3.3. As a first step, we prove in Lemma D.2 that the expected loss function $L(W)$ in Equation (7) is strictly convex with respect to (w.r.t.) $W$ and is strongly convex in any compact set of $\mathbb{R}^{d\times d}$. Moreover, we prove $L(W)$ has one unique global minimizer $W^{*}$. Since the loss function $L(W)$ we consider is highly non-linear, we cannot directly find the closed-form expression of $W^{*}$, as is often done in the prior literature. This poses a significant challenge to our analysis.
We address this technical challenge via the following method. First, in Lemma D.3, by analyzing the Taylor expansion of $L(W)$, we prove that as $N\to\infty$, our loss function $L(W)$ converges to $\widetilde{L}(W)$ pointwisely (defined in Equation (25)), and the global minimizer $W^{*}$ converges to $2\Lambda^{-1}$. Thus, we denote $W^{*}=2(\Lambda^{-1}+G)$, and prove $\|G\|_{\max}$ is bounded and scales as $\|G\|_{\max}=O(N^{-1/2})$. Next, in Lemma D.4, by further analyzing the Taylor expansion of the equation $\nabla L(W^{*})=0$ at the point $2\Lambda^{-1}$, we establish a tighter bound $\|G\|_{\max}=O(N^{-1})$. In Lemma D.5, we prove that our loss function is $l$-smooth and provide an upper bound for $l$. Thus, in a compact set $R_{W}$, our loss function is $\alpha$-strongly convex and $l$-smooth. Finally, leveraging the standard results from the convex optimization, we prove Theorem 3.3.
@@
Corollary 3.4.
If we optimize the expected loss $L(W)$ in Equation (7) via gradient descent with training prompt length $N$, initial parameters $W^{0}$, and learning rate $\eta=1/l$, then, under Assumption 3.2, after $T\geq 2\kappa\log(N\|W^{0}-W^{*}\|_{F})$ steps, the updated model $\widehat{W}$ satisfies
@@ 3.2 In-context inference
Next, we analyze the performance of the trained transformer in Equation (11) for in-context binary classification tasks. We make the following assumption.
Assumption 3.5.
@@
$\mathbb{P}\left(y_{\mathsf{query}}=1|x_{\mathsf{query}}\right)$
$=$ $\frac{f(x_{\mathsf{query}}|y_{\mathsf{query}}=1)\mathbb{P}\left(y_{\mathsf{query}}=1\right)}{\sum_{z\in\{\pm 1\}}f(x_{\mathsf{query}}|y_{\mathsf{query}}=z)\mathbb{P}\left(y_{\mathsf{query}}=z\right)}$
$=$ $\sigma((\mu_{1}-\mu_{0})^{\top}\Lambda^{-1}x_{\mathsf{query}}).$
If we test the trained transformer with parameters $\widehat{W}$ in Equation (11) and $P_{\mathsf{test}}$, by a simple calculation, we have
@@
Theorem 3.6.
Consider a test prompt $P_{\mathsf{test}}$ satisfying Assumption 3.5, and let $y_{\mathsf{query}}\sim\mathcal{P}^{b}_{y|x_{\mathsf{query}}}(\mu_{0},\mu_{1},\Lambda)$. Let $\widehat{y}_{\mathsf{query}}$ be the prediction of the trained transformer with parameters $\widehat{W}$ in Equation (11). Then, for the inference error defined in Equation (3), we have
@@
$\leq\sigma^{\prime}(\mu^{\top}\Lambda^{-1}q)\Bigg{[}\|\widehat{G}\|_{\max}\sum_{i,j\in[d]}|\mu_{i}q_{j}|$
$+\frac{1}{\sqrt{M}}\left(\frac{1}{2}|u^{\top}\Lambda^{-1}q|+\frac{2\sqrt{2}}{\sqrt{\pi}}\sum_{i,j\in[d]}|\Lambda^{-1/2}_{ij}q_{j}|\right)\Bigg{]}$
$\quad+o\left(\frac{1}{N}+\frac{1}{\sqrt{M}}\right),$
@@
The proof of Theorem 3.6 can be found in Appendix E. Since $\|\widehat{G}\|_{\max}=O(1/N)$, Theorem 3.6 suggests that if we ignore the constants regarding $\mu_{0},\mu_{1},\Lambda,x_{\mathsf{query}}$, the expected total variation distance between $y_{\mathsf{query}}$ and $\widehat{y}_{\mathsf{query}}$ is at most $O(1/N+1/\sqrt{M})$. On the other hand, for data pair $(x,y)\sim\mathcal{P}^{b}(\mu_{0},\mu_{1},\Lambda)$, the distribution of $y$, $\mathbb{P}\left(y=1|x\right)=\sigma((\mu_{1}-\mu_{0})^{\top}\Lambda^{-1}x)$, can be characterized by a logistic regression model $\sigma(w^{\top}x+b)$ with parameters $w=\Lambda^{-1}(\mu_{1}-\mu_{0})$ and $b=0$. Therefore, when $N,M\to\infty$, the prediction of the trained transformer is equivalent to the optimal logistic regressor for binary classification problems with distribution $\mathcal{P}^{b}(\mu_{0},\mu_{1},\Lambda)$.
Note that different from Assumption 3.2 which states that $\mu_{\tau,0},\mu_{\tau,1},x_{\tau,\mathsf{query}}$ are sampled according to some specific distributions during training, Assumption 3.5 does not impose strong distributional constraints on $\mu_{0},\mu_{1}$ and $x_{\mathsf{query}}$, which shows the strong generalization ability of the trained transformer. Moreover, even if $M\to\infty$, the distribution variation between $y_{\mathsf{query}}$ and $\widehat{y}_{\mathsf{query}}$ does not disappear unless $N\to\infty$. Thus, the ICL ability of trained transformers for binary classification is limited by the finite length of training prompts. Similar behaviors have also been observed in Zhang et al. (2023a) for in-context linear regression.
Remark 3.7.
Theorem 3.6 requires Assumption 3.5 to hold. For example, we need the covariance matrix $\Lambda$ in training and testing to be the same. A similar consistency requirement of $\Lambda$ in training and testing had also been observed for in-context linear regression in Zhang et al. (2023a). Here, we discuss the consequences when Assumption 3.5 does not hold. For example, suppose the labels of our data in test prompts are not balanced where $\mathbb{P}\left(y=1\right)=p_{1},\mathbb{P}\left(y=-1\right)=p_{0}$. Besides, $\mu_{0},\mu_{1}$ do not have the same $\Lambda^{-1}$ weighted norm, and the covariance matrix of test data satisfies $\Gamma\neq\Lambda$. Then, as $N,M\to\infty$, we have
$\frac{2}{M}\sum_{i=1}^{M}y_{i}x_{i}^{\top}\to 2(p_{1}\mu_{1}-p_{0}\mu_{0})^{\top},$
and
On the other hand, the distribution of the ground truth label is $\mathbb{P}\left(y_{\mathsf{query}}=1\right)=\sigma((\mu_{1}-\mu_{0})^{\top}\Gamma^{-1}x_{\mathsf{query}}+(\mu_{0}^{\top}\Lambda^{-1}\mu_{0}-\mu_{1}^{\top}\Lambda^{-1}\mu_{1})/2+\log(p_{1}/p_{0})).$ Define $z\triangleq(\mu_{1}-\mu_{0})^{\top}\Gamma^{-1}x_{\mathsf{query}}+(\mu_{1}^{\top}\Lambda^{-1}\mu_{1}-\mu_{0}^{\top}\Lambda^{-1}\mu_{0})/2+\log(p_{1}/p_{0})$ and $\hat{z}\triangleq 2(p_{1}\mu_{1}-p_{0}\mu_{0})^{\top}\Lambda^{-1}x_{\mathsf{query}}$. Then, we can see that unless $\hat{z}=z$ or $\left|\sigma(\hat{z})-\sigma(z)\right|$ is sufficiently small, the transformer cannot correctly perform the in-context binary classification.
Remark 3.8.
Another important insight of our analysis is that the pre-trained single-layer transformer can be viewed as approximately implementing linear discriminant analysis (LDA). For example, suppose we are given $\{x_{i},y_{i}\}_{i=1}^{M}$ and $x_{\mathsf{query}}$, and we need to predict the label $y_{\mathsf{query}}$ for $x_{\mathsf{query}}$. LDA assumes that $\{x_{i},y_{i}\}_{i=1}^{M}$ and $\{x_{\mathsf{query}},y_{\mathsf{query}}\}$ are i.i.d. samples, with $\mathbb{P}\left(y_{i}=1\right)=\mathbb{P}\left(y_{i}=-1\right)$, and the conditional probability density functions $f(x_{i}|y_{i}=1)$ and $f(x_{i}|y_{i}=-1)$ are Gaussian with means $\mu_{1}$, $\mu_{-1}$ and same covariance $\Sigma$. Under these assumptions, it can be derived that the optimal decision criterion for $x_{\mathsf{query}}$ is to predict $y_{\mathsf{query}}=1$ if $(\mu_{1}-\mu_{-1})^{\top}\Sigma^{-1}x_{\mathsf{query}}+\frac{1}{2}(\mu_{-1}^{\top}\Sigma^{-1}\mu_{-1}-\mu_{1}^{\top}\Sigma^{-1}\mu_{1})>0$ and $y_{\mathsf{query}}=-1$, otherwise. LDA can estimate $\hat{\mu}_{1}$ as the average of $x_{i}$ with $y_{i}=1$, estimate $\hat{\mu}_{-1}$ as the average of $x_{i}$ with $y_{i}=-1$, and estimate the covariance $\hat{\Sigma}$ from the within-class variances. For the single-layer transformer, it can compute the in-context estimate $\hat{\mu}_{1}-\hat{\mu}_{-1}=\frac{1}{M}\sum_{i=1}^{M}y_{i}x_{i}$, however, it is hard for the single-layer transformer to estimate $\hat{\Sigma}$ in context. Thus, in our paper, we make the following assumptions (Assumptions 3.2 and 3.5). We assume the pre-train data and test data have the same covariance matrix $\Lambda$ so that the transformer can learn an approximation of $\Lambda$ during pre-training. Moreover, we assume the two class means $\mu_{0}$, $\mu_{1}$ have the same $\Lambda$-weighted norm so that $\mu_{0}^{\top}\Sigma^{-1}\mu_{0}-\mu_{1}^{\top}\Sigma^{-1}\mu_{1}=0$. Under these assumptions, the quadratic term cancels out, and the estimated decision criterion simplifies to $(\frac{1}{M}\sum_{i=1}^{M}y_{i}x_{i})^{\top}\Lambda^{-1}x_{\mathsf{query}}$, which is very close to Equation (12) in our paper. When we use $\widehat{W}$ to approximate $2\Lambda^{-1}$, the estimated decision criterion becomes exactly Equation (12). Therefore, when $\widehat{W}=2\Lambda^{-1}$ and the in-context examples are balanced across classes, the transformer’s decision criterion is the same as that of the LDA with exact knowledge of $\Lambda$. Our experiments also corroborate this theoretical findings. For example, in Figure 3, since the pre-trained transformer has already learned a relatively good approximation of $\Lambda^{-1}$, while LDA must estimate $\Lambda^{-1}$ in context, the trained transformer significantly outperforms LDA when the number of in-context examples is small. As the context length increases, LDA’s performance approaches that of the trained transformer.